OpenStudy (anonymous):

Write the equation of the hyperbola with the given vertices (-1, 0) and (1,0) and with asymptotes at y = +/-3x

OpenStudy (anonymous):

@Vandreigan

OpenStudy (anonymous):

Oof, going to make me reach way back for this one. Give me a moment to remember :)

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

Alright, so first we should find the center of the hyperbola. This is the midpoint between the vertices. Do you know how to find that?

OpenStudy (anonymous):

Center is (0,0)

OpenStudy (anonymous):

Excellent :) Now, we know that the slope of the asymptotes is given by:

OpenStudy (anonymous):

\[m = \pm\frac{b}{a}\] Where a is half the distance between the vertices. Using this, find a and b.

OpenStudy (anonymous):

a = 1 b =

OpenStudy (anonymous):

My brain cant function right now haha

OpenStudy (anonymous):

How do i find b

OpenStudy (anonymous):

Plug in what you know to the equation for the slope of the asymptote. We know the slope, and we know a. Solve for b.

OpenStudy (anonymous):

+/-3x =+/1 b/1 ??

OpenStudy (anonymous):

+/-3x =+/-1 b/1 ??*

OpenStudy (anonymous):

\[\pm 3 = \pm \frac{b}{1}\] Which simplifies to: \[3 = b\]

OpenStudy (anonymous):

ohhh okay so b^2 = 9

OpenStudy (anonymous):

So, we know a, we know b, and we know where the center is. Do we know which way this hyperbola will open?

OpenStudy (anonymous):

You mean if its on the tranverse or conjugate?

OpenStudy (anonymous):

Its on transverse

OpenStudy (anonymous):

x^2/1 - y^2/9 = 1?

OpenStudy (anonymous):

Ok, so we know what equation to plug everything in to. You're ahead of me. Well done :)

OpenStudy (anonymous):

Awesome thanks for all the help :) youre a good teacher!

OpenStudy (anonymous):

My pleasure :)